OpenStudy (anonymous):

Solve for x:

OpenStudy (anonymous):

\[\frac{ 5 }{ x^2-1 } - \frac{ 2 }{ x } = \frac{ 2 }{ x+1 }\]

OpenStudy (anonymous):

\[\frac{ 5 }{ (x-1)(x+1) }-\frac{ 2 }{ x }=\frac{ 2 }{ x+1 }\] \[5x-2((x-1)(x+1))=2(x-1)(x)\] 5x-2x^2+2=2x^2-2x \[0=4x ^{2}-7x-2\]

OpenStudy (anonymous):

you finish with the quadratic formula

OpenStudy (anonymous):

what is the quadratic formula?

OpenStudy (anonymous):

the answer to the quadratic formula will give you 1 or 2 values for x (it can also give you an imaginary solution if you have a negative in your sqrt, but that's something you'll see later)

OpenStudy (anonymous):

\[ax^{2}+bx+c=0\] quadratic equation to find x is : \[\frac{ -b }{ 2a } \pm \frac{ \sqrt{b ^{2}-4ac} }{ 2a }\]